THE ROLE OF A NOVEL DIRADICAL PATHWAY IN REACTIONS BETWEEN PEROXYL RADICALS AND NO

\underline{EVAN B. JOCHNOWITZ}, S. J. BLANKSBY, G. B. ELLISON, Dept Chemistry, Univ. Colo. Boulder, CO 80309-0215; J. F. STANTON, Dept. Chemistry, University of Texas, Austin, TX 78712.

The conversion of peroxyl radicals to organic nitrates via reaction with NO is of importance in atmospheric chemistry and biochemistry. The mechanism for nitrate formation is obscure; no previous theoretical results have been even vaguely consistent with the experimental evidence. We propose a simple valence bond argument to rationalize how an initially formed pernitrite, ROONO, can decompose to an alkoxy radical and NO_{2} or rearrange to RONO_{2}. This qualitative mechanism, which involves the coupling of two valence bond states, is supported by coupled-cluster electronic structure calculations that predict a barrier of ca. 20-30 kcal mol^{−1}, that is provided by the radical/radical adduct. In addition to its obvious importance for atmospheric chemistry, it is likely that this mechanism is also responsible for chain termination of the oxygen-dependent oxidation of lipid membranes in biological cells. The mechanism can likewise be applied to the thermal decomposition of organic nitro compounds (explosives and solid propellants), RNO_{2} + heat → products.